# Beginner’s Charm: Beginner-Heavy Teams Are Associated With High Scientific Disruption

### v1

@@ Discussion
All Data 28,042,889 0.1062∗∗∗ 0.0810∗∗∗ 0.0621∗∗∗
Physics 3,292,761 0.1242∗∗∗ 0.1060∗∗∗ 0.0821∗∗∗
Internal medicine 2,309,009 0.0709∗∗∗ 0.0451∗∗∗ 0.0339∗∗∗
Statistics 1,620,324 0.1205∗∗∗ 0.1150∗∗∗ 0.0894∗∗∗
Optics 1,429,519 0.0844∗∗∗ 0.0695∗∗∗ 0.0537∗∗∗
Engineering 1,254,565 0.1129∗∗∗ 0.0926∗∗∗ 0.0707∗∗∗
Operating system 1,077,678 0.1285∗∗∗ 0.1257∗∗∗ 0.0964∗∗∗
Finance 844,720 0.1372∗∗∗ 0.1329∗∗∗ 0.1038∗∗∗
Electrical engineering 833,686 0.1103∗∗∗ 0.1012∗∗∗ 0.0773∗∗∗
Epistemology 794,365 0.0980∗∗∗ 0.0921∗∗∗ 0.0723∗∗∗
Algorithm 776,134 0.1112∗∗∗ 0.1040∗∗∗ 0.0813∗∗∗
Philosophy 765,363 0.0804∗∗∗ 0.0553∗∗∗ 0.0430∗∗∗
Psychiatry 651,434 0.0813∗∗∗ 0.0481∗∗∗ 0.0361∗∗∗
Anatomy 611,044 0.0610∗∗∗ 0.0410∗∗∗ 0.0313∗∗∗
Oceanography 529,146 0.1128∗∗∗ 0.0991∗∗∗ 0.0765∗∗∗
Organic chemistry 484,949 0.0830∗∗∗ 0.0680∗∗∗ 0.0517∗∗∗
Chemistry 482,613 0.0954∗∗∗ 0.0723∗∗∗ 0.0548∗∗∗
Sociology 458,611 0.0956∗∗∗ 0.0525∗∗∗ 0.0408∗∗∗
Combinatorics 384,339 0.0909∗∗∗ 0.0945∗∗∗ 0.0747∗∗∗
Geology 367,807 0.1423∗∗∗ 0.1323∗∗∗ 0.1020∗∗∗
Environmental chemistry 341,056 0.1246∗∗∗ 0.1071∗∗∗ 0.0818∗∗∗
Art 323,605 0.0553∗∗∗ 0.0291∗∗∗ 0.0229∗∗∗
Pathology 323,486 0.0803∗∗∗ 0.0496∗∗∗ 0.0372∗∗∗
Machine learning 302,811 0.1198∗∗∗ 0.1152∗∗∗ 0.0898∗∗∗
Surgery 274,228 0.0588∗∗∗ 0.0358∗∗∗ 0.0268∗∗∗
Ecology 251,538 0.0823∗∗∗ 0.0598∗∗∗ 0.0462∗∗∗
Composite material 242,925 0.0927∗∗∗ 0.0794∗∗∗ 0.0604∗∗∗
Genetics 237,670 0.0623∗∗∗ 0.0269∗∗∗ 0.0203∗∗∗
Horticulture 236,497 0.1037∗∗∗ 0.0924∗∗∗ 0.0704∗∗∗
Psychology 223,317 0.1103∗∗∗ 0.0776∗∗∗ 0.0596∗∗∗
Aesthetics 215,875 0.0702∗∗∗ 0.0657∗∗∗ 0.0521∗∗∗
Human–computer interaction 201,350 0.1151∗∗∗ 0.1129∗∗∗ 0.0863∗∗∗
Endocrinology 191,578 0.0856∗∗∗ 0.0626∗∗∗ 0.0474∗∗∗
Mathematical analysis 187,048 0.0967∗∗∗ 0.0920∗∗∗ 0.0721∗∗∗
Biophysics 180,521 0.0408∗∗∗ 0.0303∗∗∗ 0.0233∗∗∗
Business 180,104 0.1044∗∗∗ 0.1033∗∗∗ 0.0803∗∗∗
Gender studies 166,773 0.0821∗∗∗ 0.0675∗∗∗ 0.0535∗∗∗
Developmental psychology 164,434 0.1059∗∗∗ 0.0743∗∗∗ 0.0570∗∗∗
World Wide Web 160,372 0.0920∗∗∗ 0.0549∗∗∗ 0.0421∗∗∗
Artificial intelligence 159,812 0.1256∗∗∗ 0.1177∗∗∗ 0.0908∗∗∗
Virology 129,810 0.0943∗∗∗ 0.0484∗∗∗ 0.0363∗∗∗
Physical chemistry 124,339 0.0822∗∗∗ 0.0682∗∗∗ 0.0522∗∗∗
Animal science 118,178 0.1100∗∗∗ 0.0943∗∗∗ 0.0716∗∗∗
Psychoanalysis 114,169 0.0732∗∗∗ 0.0454∗∗∗ 0.0350∗∗∗
Geomorphology 112,980 0.1446∗∗∗ 0.1343∗∗∗ 0.1034∗∗∗
Particle physics 112,792 0.0689∗∗∗ 0.0491∗∗∗ 0.0388∗∗∗
Nuclear chemistry 109,607 0.0904∗∗∗ 0.0753∗∗∗ 0.0570∗∗∗
Ophthalmology 105,613 0.0735∗∗∗ 0.0413∗∗∗ 0.0313∗∗∗
Environmental resource management 100,286 0.1343∗∗∗ 0.1236∗∗∗ 0.0955∗∗∗
Paleontology 92,851 0.0848∗∗∗ 0.0500∗∗∗ 0.0383∗∗∗
Literature 91,376 0.0532∗∗∗ 0.0337∗∗∗ 0.0265∗∗∗
Nuclear engineering 91,294 0.0907∗∗∗ 0.0773∗∗∗ 0.0592∗∗∗
Advertising 85,553 0.1205∗∗∗ 0.1102∗∗∗ 0.0859∗∗∗
Physical geography 82,795 0.1386∗∗∗ 0.1181∗∗∗ 0.0918∗∗∗
Public economics 80,439 0.1051∗∗∗ 0.1080∗∗∗ 0.0860∗∗∗
Financial economics 74,294 0.0757∗∗∗ 0.0684∗∗∗ 0.0528∗∗∗
Mathematics education 73,267 0.0963∗∗∗ 0.0963∗∗∗ 0.0743∗∗∗
Distributed computing 71,372 0.0880∗∗∗ 0.0872∗∗∗ 0.0675∗∗∗
Immunology 70,808 0.0827∗∗∗ 0.0337∗∗∗ 0.0252∗∗∗
Software engineering 68,734 0.1082∗∗∗ 0.0987∗∗∗ 0.0759∗∗∗
Biochemistry 66,719 0.0660∗∗∗ 0.0347∗∗∗ 0.0265∗∗∗
Data mining 66,278 0.1204∗∗∗ 0.1180∗∗∗ 0.0909∗∗∗
Computer science 65,752 0.0474∗∗∗ 0.0585∗∗∗ 0.0447∗∗∗
Internet privacy 62,716 0.1057∗∗∗ 0.0916∗∗∗ 0.0711∗∗∗
Biotechnology 56,711 0.0958∗∗∗ 0.0692∗∗∗ 0.0523∗∗∗
@@
8+ -0.964 [-1.000, -0.829]*** -0.200 [-0.796, 0.520] 0.491 [-0.208, 0.956]
Table S5: Disruption percentiles by author-ratio bins. Cells show median disruption percentile and sample size ($n$) for the left subplot of Figure 3: Early Career Author Ratio (rows) vs Beginner Author Ratio (columns). .
Beginner Author Ratio (bin center)
Early Career Ratio 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
1.0 56.45 n=2.95M
0.9 47.40 n=33.07K 59.38 n=8.79K
0.8 48.06 n=390.09K 49.06 n=3.72K 59.38 n=92.63K
0.7 46.76 n=1.10M 49.38 n=42.69K 50.71 n=1.40K 63.44 n=234.27K
0.6 42.96 n=337.57K 44.04 n=43.26K 51.75 n=98.10K 52.13 n=9.59K 63.42 n=24.49K
0.5 45.61 n=3.61M 44.94 n=104.07K 52.79 n=281.42K 54.76 n=28.49K 55.55 n=3.30K 63.24 n=581.68K
0.4 40.68 n=632.59K 40.97 n=105.63K 47.29 n=221.65K 47.59 n=24.70K 59.38 n=81.78K 59.38 n=2.87K 65.50 n=21.60K
0.3 43.64 n=2.14M 42.08 n=226.95K 42.15 n=16.97K 53.59 n=750.14K 52.35 n=16.23K 59.38 n=20.29K 59.38 n=6.29K 67.42 n=145.50K
0.2 42.17 n=1.45M 39.09 n=72.95K 47.85 n=690.19K 44.62 n=8.98K 52.73 n=107.93K 59.38 n=151.90K 59.38 n=40.79K 63.27 n=1.24K 68.32 n=55.40K
0.1 38.28 n=424.30K 41.39 n=230.54K 41.61 n=26.76K 46.92 n=87.59K 48.41 n=23.08K 55.66 n=28.91K 59.38 n=10.12K 59.38 n=16.17K 64.23 n=2.21K 67.08 n=7.70K
0.0 51.40 n=6.77M 41.63 n=102.27K 47.91 n=389.50K 51.98 n=593.19K 51.44 n=65.81K 55.78 n=1.06M 59.38 n=31.36K 59.38 n=196.12K 63.97 n=61.31K 64.09 n=7.85K 70.25 n=928.71K
@@
Beginner Author Ratio (bin center)
Senior Author Ratio 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
1.0 51.44 n=6.75M
0.9 38.33 n=415.52K 41.78 n=98.23K
0.8 42.18 n=1.45M 38.54 n=51.09K 47.94 n=387.47K
0.7 43.62 n=2.15M 41.58 n=211.59K 40.93 n=11.84K 51.99 n=592.06K
0.6 40.61 n=639.06K 40.00 n=138.00K 45.33 n=259.66K 44.85 n=32.45K 51.47 n=65.14K
0.5 45.58 n=3.62M 42.73 n=172.17K 48.97 n=452.45K 47.84 n=58.64K 45.90 n=8.27K 55.78 n=1.06M
0.4 42.86 n=342.50K 41.50 n=90.47K 47.14 n=225.43K 45.48 n=38.27K 52.44 n=119.89K 51.94 n=4.46K 59.38 n=30.96K
0.3 46.74 n=1.10M 45.45 n=114.03K 43.56 n=10.33K 53.77 n=739.61K 51.79 n=16.93K 56.04 n=25.71K 55.64 n=7.87K 59.38 n=195.78K
0.2 48.03 n=391.45K 44.25 n=13.52K 52.59 n=373.88K 48.07 n=4.51K 59.38 n=76.33K 59.38 n=151.75K 59.38 n=41.12K 59.38 n=1.35K 63.98 n=61.00K
0.1 47.31 n=33.63K 49.92 n=42.79K 51.16 n=4.68K 54.30 n=36.87K 55.90 n=11.27K 59.38 n=22.25K 59.38 n=8.30K 59.38 n=16.09K 63.62 n=2.19K 63.86 n=7.56K
0.0 56.45 n=2.95M 59.38 n=8.99K 59.38 n=92.87K 63.44 n=234.53K 63.39 n=24.77K 63.24 n=581.98K 65.40 n=21.91K 67.41 n=145.81K 68.30 n=55.72K 67.28 n=7.99K 70.25 n=928.71K
@@
Early Career Author Ratio (bin center)
Senior Author Ratio 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
1.0 51.44 n=6.75M
0.9 40.98 n=109.13K 38.42 n=404.62K
0.8 47.77 n=393.18K 37.74 n=63.35K 42.28 n=1.43M
0.7 51.92 n=594.87K 42.07 n=194.95K 37.21 n=46.92K 43.71 n=2.12M
0.6 51.25 n=66.65K 43.73 n=49.12K 44.41 n=289.08K 39.56 n=109.67K 40.75 n=619.79K
0.5 55.77 n=1.06M 47.60 n=64.27K 49.03 n=450.63K 43.40 n=139.73K 39.33 n=50.37K 45.62 n=3.60M
0.4 59.38 n=31.53K 50.51 n=19.68K 52.22 n=111.34K 44.72 n=44.15K 45.80 n=281.11K 40.94 n=30.07K 42.98 n=334.09K
0.3 59.38 n=196.23K 56.08 n=32.50K 48.95 n=4.48K 53.93 n=732.14K 46.39 n=29.56K 46.03 n=83.02K 43.65 n=36.67K 46.77 n=1.10M
0.2 63.95 n=61.38K 59.38 n=3.71K 59.38 n=189.71K 51.76 n=3.47K 56.20 n=77.24K 52.85 n=280.27K 50.95 n=104.68K 44.10 n=4.85K 48.06 n=389.58K
0.1 63.81 n=7.87K 62.49 n=17.14K 59.38 n=3.28K 59.38 n=25.69K 59.38 n=10.85K 54.99 n=30.62K 51.64 n=12.82K 50.07 n=40.41K 48.79 n=4.03K 47.38 n=32.93K
0.0 70.24 n=929.10K 67.13 n=8.03K 68.27 n=55.73K 67.41 n=145.82K 65.42 n=21.89K 63.24 n=581.95K 63.39 n=24.76K 63.43 n=234.50K 59.38 n=92.82K 59.38 n=8.93K 56.45 n=2.95M

### v2

@@ 4.3 Atypical Combination Score
All Data 29,054,261 0.107232*** 0.081831*** 0.062813***
1941-1950 47,213 0.042378*** 0.048536*** 0.038080***
1951-1960 235,753 0.046042*** 0.043221*** 0.033310***
1961-1970 730,667 0.072997*** 0.058709*** 0.045113***
1971-1980 1,477,240 0.068561*** 0.053265*** 0.041002***
1981-1990 2,418,102 0.073413*** 0.052354*** 0.040219***
1991-2000 3,757,344 0.075212*** 0.051006*** 0.038985***
2001-2010 7,290,170 0.096032*** 0.067303*** 0.051502***
2011-2020 13,097,772 0.137403*** 0.109909*** 0.084223***
Table S3: Correlations between author beginner ratio and disruption percentile by discipline in the overall dataset. Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$ with significance stars: $^{***}P<0.001$.
Group NN Pearson $r$ Spearman $\rho$ Kendall $\tau$
Art 1,106,563 0.0850∗∗∗ 0.0751∗∗∗ 0.0585∗∗∗
Biology 8,240,399 0.0972∗∗∗ 0.0679∗∗∗ 0.0519∗∗∗
Business 1,721,427 0.1263∗∗∗ 0.1255∗∗∗ 0.0974∗∗∗
Chemistry 7,740,170 0.0976∗∗∗ 0.0797∗∗∗ 0.0611∗∗∗
Computer science 9,102,699 0.1193∗∗∗ 0.1118∗∗∗ 0.0865∗∗∗
Economics 2,670,974 0.1196∗∗∗ 0.1115∗∗∗ 0.0870∗∗∗
Engineering 5,735,626 0.1162∗∗∗ 0.1048∗∗∗ 0.0804∗∗∗
Environmental science 1,418,544 0.1424∗∗∗ 0.1297∗∗∗ 0.0995∗∗∗
Geography 1,710,299 0.1248∗∗∗ 0.1115∗∗∗ 0.0862∗∗∗
Geology 1,875,846 0.1351∗∗∗ 0.1218∗∗∗ 0.0943∗∗∗
History 713,100 0.0751∗∗∗ 0.0573∗∗∗ 0.0450∗∗∗
Materials science 5,159,780 0.1025∗∗∗ 0.0883∗∗∗ 0.0680∗∗∗
Mathematics 5,968,273 0.1184∗∗∗ 0.1140∗∗∗ 0.0891∗∗∗
Medicine 7,501,956 0.0867∗∗∗ 0.0558∗∗∗ 0.0421∗∗∗
Philosophy 3,129,726 0.1066∗∗∗ 0.0943∗∗∗ 0.0736∗∗∗
Physics 8,476,490 0.1141∗∗∗ 0.0999∗∗∗ 0.0774∗∗∗
Political science 2,898,145 0.1007∗∗∗ 0.0917∗∗∗ 0.0720∗∗∗
Psychology 3,472,859 0.1119∗∗∗ 0.0872∗∗∗ 0.0672∗∗∗
Sociology 2,462,381 0.0983∗∗∗ 0.0747∗∗∗ 0.0582∗∗∗
@@
8+ -0.970 [-1.000, -0.811]*** -0.333 [-0.806, 0.310] 0.485 [-0.153, 0.931]
Table S5: Disruption percentiles by author-ratio bins. Cells show median disruption percentile and sample size ($n$) for the left subplot of Figure 3: Early Career Author Ratio (rows) vs Beginner Author Ratio (columns).
Early Career Ratio Beginner Author Ratio (bin center)
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
1.0 58.92 n=3.34M
0.9 47.07 n=33.60K 58.92 n=9.00K
0.8 47.79 n=396.08K 48.60 n=3.85K 58.92 n=96.58K
0.7 46.55 n=1.12M 49.06 n=43.26K 50.63 n=1.52K 62.87 n=251.87K
0.6 42.81 n=340.80K 43.87 n=43.83K 51.37 n=99.33K 51.86 n=9.75K 62.96 n=25.23K
0.5 45.49 n=3.69M 44.80 n=103.52K 52.44 n=286.04K 54.32 n=28.60K 54.95 n=3.28K 62.80 n=657.12K
0.4 40.61 n=633.04K 40.89 n=105.69K 47.06 n=222.45K 47.35 n=24.77K 58.92 n=82.69K 58.92 n=2.85K 65.03 n=22.20K
0.3 43.51 n=2.15M 41.98 n=227.08K 42.05 n=16.98K 53.26 n=770.09K 52.00 n=16.27K 58.92 n=20.31K 58.92 n=6.43K 66.95 n=157.57K
0.2 42.06 n=1.45M 39.05 n=72.95K 47.61 n=693.13K 44.43 n=8.98K 52.38 n=108.40K 58.92 n=154.84K 58.92 n=41.41K 58.92 n=1.30K 67.96 n=57.95K
0.1 38.26 n=424.27K 41.31 n=230.57K 41.52 n=26.76K 46.70 n=87.66K 48.16 n=23.10K 55.25 n=28.70K 58.92 n=10.24K 58.92 n=16.39K 63.68 n=2.26K 66.53 n=8.00K
0.0 51.26 n=6.90M 41.54 n=102.26K 47.66 n=389.98K 51.67 n=598.23K 51.13 n=65.92K 55.51 n=1.11M 58.92 n=31.56K 58.92 n=202.78K 63.50 n=62.56K 63.45 n=7.95K 70.37 n=1.10M
@@
Senior Author Ratio Beginner Author Ratio (bin center)
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
1.0 51.28 n=6.89M
0.9 38.28 n=420.01K 41.68 n=98.80K
0.8 42.07 n=1.45M 38.52 n=52.13K 47.69 n=388.24K
0.7 43.49 n=2.16M 41.49 n=212.58K 41.05 n=12.39K 51.68 n=597.27K
0.6 40.58 n=636.34K 39.94 n=138.23K 45.13 n=260.45K 44.72 n=32.75K 51.15 n=65.34K
0.5 45.48 n=3.69M 42.67 n=169.50K 48.74 n=453.76K 47.59 n=58.25K 45.61 n=8.19K 55.51 n=1.11M
0.4 42.75 n=343.40K 41.40 n=90.53K 46.92 n=226.22K 45.30 n=38.30K 52.09 n=120.37K 51.59 n=4.38K 58.92 n=31.20K
0.3 46.53 n=1.12M 45.28 n=114.33K 43.44 n=10.34K 53.44 n=759.59K 51.46 n=16.98K 55.59 n=25.59K 55.22 n=8.00K 58.92 n=202.46K
0.2 47.78 n=396.88K 44.10 n=13.56K 52.24 n=379.84K 47.81 n=4.52K 58.92 n=77.20K 58.92 n=154.64K 58.92 n=41.77K 58.92 n=1.44K 63.50 n=62.31K
0.1 47.02 n=33.92K 49.62 n=43.17K 50.89 n=4.72K 53.86 n=37.16K 55.55 n=11.35K 58.92 n=22.18K 58.92 n=8.42K 61.97 n=16.32K 63.19 n=2.25K 63.38 n=7.71K
0.0 58.92 n=3.34M 58.92 n=9.18K 58.92 n=96.80K 62.87 n=252.11K 62.93 n=25.46K 62.80 n=657.25K 64.96 n=22.48K 66.93 n=157.84K 67.95 n=58.22K 66.47 n=8.24K 70.37 n=1.10M
@@
Sen. Auth. Ratio Early Career Author Ratio (bin center)
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
1.0 51.28 n=6.89M
0.9 41.19 n=105.89K 38.33 n=412.91K
0.8 47.60 n=391.71K 38.03 n=60.02K 42.13 n=1.44M
0.7 51.64 n=599.01K 42.11 n=192.55K 37.49 n=43.62K 43.55 n=2.14M
0.6 51.07 n=66.31K 43.75 n=47.90K 44.39 n=286.62K 39.67 n=107.06K 40.64 n=625.24K
0.5 55.51 n=1.11M 47.53 n=63.00K 48.84 n=450.49K 43.45 n=135.26K 39.65 n=45.38K 45.50 n=3.69M
0.4 58.92 n=31.69K 50.21 n=19.70K 51.88 n=111.81K 44.54 n=44.19K 45.61 n=281.95K 41.06 n=27.20K 42.82 n=337.88K
0.3 58.92 n=202.82K 55.64 n=32.60K 48.68 n=4.48K 53.60 n=752.09K 46.17 n=29.63K 45.95 n=81.63K 43.51 n=36.95K 46.55 n=1.12M
0.2 63.48 n=62.63K 58.92 n=3.72K 58.92 n=193.39K 51.40 n=3.48K 55.75 n=78.11K 52.53 n=284.47K 50.61 n=105.71K 43.98 n=5.05K 47.79 n=395.59K
0.1 63.35 n=7.96K 62.15 n=17.28K 58.92 n=3.29K 58.92 n=25.90K 58.92 n=10.92K 54.63 n=30.52K 51.44 n=12.93K 49.77 n=40.80K 48.08 n=4.16K 47.07 n=33.43K
0.0 70.36 n=1.10M 66.60 n=8.26K 67.90 n=58.22K 66.93 n=157.84K 64.96 n=22.46K 62.80 n=657.23K 62.95 n=25.47K 62.86 n=252.09K 58.92 n=96.76K 58.92 n=9.17K 58.92 n=3.34M
Table S8: Kendall’s $\tau$ between author ratio and disruption percentile by team size and career stage (for the 4-way career-stage split case). Entries are Kendall’s $\tau$ with bootstrap 95% percentile CIs in brackets; significance stars reflect adjusted $P$-values: $^{*}P<0.05$, $^{**}P<0.01$, $^{***}P<0.001$. Benjamini–Hochberg FDR correction applied across all tests.
Team size Beginner Early-career Mid-career Senior
1 1.000 [1.000, 1.000]
2 1.000 [1.000, 1.000] 0.333 [-1.000, 1.000] 0.333 [-1.000, 1.000] -1.000 [-1.000, -1.000]
3 1.000 [1.000, 1.000] 0.667 [-1.000, 1.000] 0.667 [-1.000, 1.000] -1.000 [-1.000, -1.000]
4 1.000 [1.000, 1.000]* 0.800 [0.000, 1.000] 0.800 [0.000, 1.000] -1.000 [-1.000, -1.000]*
5 1.000 [1.000, 1.000]** 0.867 [0.385, 1.000]* 0.867 [0.385, 1.000]* -1.000 [-1.000, -1.000]**
6 1.000 [1.000, 1.000]** 0.810 [0.250, 1.000]* 0.905 [0.556, 1.000]** -1.000 [-1.000, -1.000]**
7 1.000 [1.000, 1.000]*** 0.857 [0.410, 1.000]** 0.810 [0.294, 1.000]* -1.000 [-1.000, -1.000]***
8+ 0.931 [0.705, 1.000]*** 0.911 [0.600, 1.000]*** 0.889 [0.484, 1.000]** -0.939 [-1.000, -0.729]***